# Uncertainty-aware Reward Model: Teaching Reward Models to Know What is Unknown

### v1

@@ 4.2 Uncertainty-aware Reward Model Ensemble
Bootstrap ensemble of models is simple and effective for epistemic uncertainty quantification compared with other methods (Neal, 2012; Hernández-Lobato & Adams, 2015; Blundell et al., 2015).
Figure 3: Illustration of URME in epistemic uncertainty quantification.
Fig. 3 illustrates how URME works in quantifying uncertainty. The input space $X\times Y$ ($X$ for prompt and $Y$ for response) is split into the known and unknown area. In the known area, the training dataset can well support URMs to make reliable reward evaluations. However, in the unknown area, the situation differs. Given that each model utilizes different weight initialization and is optimized with distinct data mini-batches, the ensemble models are likely to diverge, potentially leading to varying and inconsistent evaluations. This inconsistency and divergence among models indicates the degree of uncertainty in URME and large uncertainty in turn indicates the input data may belong to the unknown area.
@@
Accurate reward evaluation is crucial in LLM alignment, as it fundamentally steers the learning process. Thus, we can adopt a filtering strategy to discard data with highly uncertain reward evaluations, since RMs may exhibit poor generalization and lack sufficient knowledge to provide reliable feedbacks for them. In this way, we can prevent LLMs from learning undesired behaviors, promoting a more controlled and trustworthy alignment process.
@@ 5 Experiment
Table 1: Results on RewardBench. RewardBench evaluates four abilities: Chat, Chat-Hard (C-HARD), Safety and Reasoning. Rankings are decided by the overall average score of all categories. Results of other models are taken from the RewardBench Leaderboard (by 23rd Sept. 2024).
MODEL BASE SCORE CHAT C-HARD SAFETY REASON
URM (Ours) Llama3.1-8B 92.9 95.5 88.2 91.1 97.0
SFR-Judge-r Llama3.1-70B 92.7 96.9 84.8 91.6 97.6
Skywork-8B Llama3.1-8B 92.5 95.8 87.3 90.8 96.2
Nemotron-RM Nemotron4-340B 92.0 95.8 87.1 91.5 93.6
GRM Llama3-8B 91.5 95.5 86.2 90.8 93.6
ArmoRM Llama3-8B 90.4 96.9 76.8 90.5 97.3
InternLM2-RM InternLM2-20B 90.2 98.9 76.5 89.5 95.8
SteerLM-RM Llama3-70B 88.8 91.3 80.3 92.8 90.6
Gemini-1.5-pro - 88.2 92.3 80.6 87.9 92.0
GPT-4o - 86.7 96.1 76.1 88.1 86.6
GPT-4-turbo - 86.0 95.3 74.3 87.6 86.9
@@ 5.1 Experiment Settings
In our experiment, URM is based on Llama3.1 with 8 billion parameters. Before adding the probabilistic value head, we initialize URM’s base model with weights from Liu & Zeng (2024). The gating layer consists of two fully-connected layers with hidden size 4096 activated by SELU (Klambauer et al., 2017). More information on URM training and implementation is given in the appendix A.1, C.1. URME have 3 URMs with different random seeds, probabilistic value head initialization and mini-batches of training data.
We utilize HelpSteer 2 (Wang et al., 2024c) as the training dataset to train the base model and the probabilistic value head for 1 epoch with learning rate $2\times 10^{-6}$. After obtaining the attribute-specific uncertain-aware probabilistic value head and base model, we keep them frozen and train the gating layer on Skywork-reward-preference-80k (Liu & Zeng, 2024) for 4000 steps with batchsize 256. During training, we held out 4k data from the dataset as validation set to choose the checkpoint with highest validation accuracy.
RewardBench (Lambert et al., 2024) , our evaluation benchmark for RMs, has 2985 questions and response pairs. For multi-attribute RMs and BT-model RMs, a prediction for a response pair is correct if the RM gives a higher reward to the chosen response than the rejected response. For generative models, RewardBench evaluates them via LLM-as-a-judge (Zheng et al., 2023). If the generative model prioritizes the chosen response than the rejected response, the prediction is seen as correct. To test URM and URME’s ability in improving LLMs’ generation quality, we evaluate URM and URME with best-of-$n$ sampling (Stiennon et al., 2020) on AlpacaEval (Li et al., 2023).
@@ 5.2.1 Overall Results
Table 1 gives the results on RewardBench. The compared baselines include multi-attribute RMs (Nemotron4-Reward (Adler et al., 2024), ArmoRM (Wang et al., 2024a), SteerLM-RM (Dong et al., 2023b)), BT-model RMs (Skywork-reward (Liu & Zeng, 2024), GRM (Yang et al., 2024), InternLM2-RM (Cai et al., 2024)) and generative RMs (SFR-Judge-r, Gemini-1.5-pro (Google, 2024), GPT-4o (OpenAI, 2024b), GPT-4-turbo (OpenAI, 2024a)). SFR-Judge-r is a chatbot developed by Salesforce based on Llama3.1-70B.
The results on RewardBench confirm URM’s strong ability in reward modeling. URM achieves the highest ranking among 8B models and outperforms a number of larger models including Nemotron-4-340B-Reward, also a multi-attribute RM. Except Chat where almost all models have relatively good performance, URM demonstrates improvement over the base model Skywork-8B in all abilities. Especially, compared to ArmoRM which is also a multi-attribute RM with gating layers, URM’s better performance shows the efficacy of modeling human preferences as distributions.
Figure 4: RewardBench overall scores for the ablation study. URM_Reg is URM trained with attribute regression while URM_MLE is trained via maximum likelihood estimation. ’Ablation’ replaces the uncertainty-aware value head with a linear layer directly predicting the attribute scores.
@@ 5.2.2 Ablation Study
Here we study the effect of the uncertain-aware value head and different training methods of URMs. To test the applicability of URM, we initialize URM with two different base models: Skywork-RM (Liu & Zeng, 2024) and Fsfairx-RM (Dong et al., 2023a). Fig. 4 gives the results of our ablation study. ’Ablation’ refers to the model with a value head to directly map hidden states to score values instead of sampling in URM. All other components of Ablation are kept the same as URM. URM_Reg is an URM trained with the attribute regression loss function in Eq. 5, while URM_MLE is trained via maximum likelihood estimation. Since the dataset Helpsteer 2 for our attribute prediction has already been used in the base model Skywork-RM, Ablation and URM_MLE do not demonstrate improvement over the base model, and only URM_Reg surpasses the base model by modeling the preference distributions. But with base model Fsfairx-RM not trained with Helpsteer 2 previously, all our models demonstrate significant improvement over the base model. Especially, URM trained via attribute regression significantly outperform its counterpart with MLE loss. However, although URM_Reg has better performance in prioritizing chosen responses over the rejected, URM_MLE demonstrates better uncertainty quantification and distribution modeling ability. We theoretically illustrate this phenomenon in the appendix B. Thus, for other studies involving uncertainty quantification, we use URMs trained via the MLE loss.
Fig. 4 indicates regression-based training methods achieve higher scores on RewardBench. This could potentially be credited to the high quality of Helpsteer 2 dataset, which is meticulously processed and derived from Helpsteer (Wang et al., 2023). This quality enables even the simplest direct attribute regression to deliver substantial performance improvements, as shown by the Ablation with base model Fsfairx-RM. However, the introduction of noise via the sampling-based scores in URM_Reg makes URMs more robust in distinguishing between chosen and rejected responses. Despite this, we anticipate that URM_MLE would prove more successful on real-world datasets, which often encompass a wide spectrum of data quality, so that modeling distributions of the scores becomes more necessary.
@@ 5.2.3 Uncertainty Quantification
Now we study the uncertainty quantification of URM and URME and how they behave when dealing in-distribution and OOD data. Given the challenge inherent in identifying what precisely is OOD for LLMs, we adopt numeric calculations as simulated OOD data. This is because LLMs are known to underperform in this skill area. Details are given in the appendix A.2.
@@
Fig. 5 gives the attribution score attributions of OOD and in-distribution data modeled by URM. Due to the lack of knowledge to accurately evaluate the OOD data, the modeled distributions for OOD data have significantly larger variance and are much closer to uniformity than for in-distributional data. Therefore, this uncertainty quantified by the variance can serve as an informative tool for identifying and filtering out OOD data, where reward models exhibit a tendency towards making uniform guess than providing an accurate evaluation. This strategy ensures the evaluated outcomes are both more dependable and robust.
Figure 5: Attribute score distributions modeled by URM. Means and variances are estimated and averaged by OOD and in-distributional samples separately.
Figure 6: URME’s uncertainty quantification by (a) maximum reward gaps and (b) maximum covariance norms. Larger discrepancies exist among the URM ensemble when dealing with OOD data. (c) Reward evaluation accuracy when uncertainty of prompt-response pairs is within the threshold. 2000 test questions are used in this evaluation. The results confirm that uncertainty of URM and URME is able to indicate reliability of reward predictions.
@@
We quantify uncertainty in URME with two metrics: maximum reward gaps in Eq. 7 and maximum covariance norms in Eq. 8. URME uncertainty quantification results are given in Fig. 6(a), (b). Quantified uncertainty under two metrics both indicate that URME is substantially more uncertain on OOD data. The results confirm that when the URMs lack relevant knowledge to make accurate reward predictions, they will diverge with each other, demonstrating significant discrepancies.
@@
To test whether quantifying uncertainty is able to improve reliability of the evaluated rewards, we utilize 2k prompts from our held-out validation set (described earlier) and evaluate their rewards and uncertainties. Fig. 6(c) gives URM and URME’s evaluation accuracy with uncertainty-based filtering. In this setup, prompts and responses (either the chosen or rejected) with uncertainty larger than the threshold are filtered out. URME use reward gaps to quantify uncertainty and URM’s uncertainty is quantified by summation of each attribute distribution’s variance. The results validate our claim that reward predictions with low uncertainty are more reliable than those with high uncertainty. Therefore, through uncertainty quantification, we can decide whether the reward predictions are unreliable and need to be filtered out, which will lead to improved reliability of reward evaluations.
5.2.4 Generation Results Improvement
We evaluate URM and URME’s ability in improving LLMs’ generations with best-of-$n$ sampling on AlpacaEval (Li et al., 2023). Specifically, we prompt Llama3-8b-Instruct model with 805 questions from AlpacaEval for $n$ times, evaluate each response’s reward and choose the response with highest reward (highest average reward for URME) as the answer. The answer is then compared against the reference answer provided by the benchmark with LLM-as-a-judge (Zheng et al., 2023). In LLM-as-a-judge, we use the official prompt of AlpacaEval and GPT-4-0125-preview as the judge. Details are given in the appendix A.3.
Table 2: Win rates of Llama3-8b-Instruct against reference answers on AlpacaEval.
Evaluator Best-of-1 Best-of-4 Best-of-8 Best-of-16 Best-of-32 Best-of-64
URM 81.2% 82.7% 83.1% 83.5% 83.9% 85.3%
URME 81.2% 83.4% 84.5% 85.5% 86.6% 86.4%
Table 2 gives the results of using URM and URME to improve generation quality. In our experiments, the baseline model Llama3-8b-Instruct achieves $81.2\%$ win rate (best-of-1). As the number of samples increases, both URM and URME are able to evaluate the quality of responses, thus ameliorating the baseline model’s generative performance. Furthermore, URME is able to consistently outperform URM in improving generation quality, as it combines the strength of several independent models and mitigates biases during reward evaluation (Coste et al., 2023; Eisenstein et al., 2023).
6 Conclusion and Future Work
In this paper, we study the uncertainty issue in reward modeling for LLMs. Uncertainty-aware Reward Model (URM) and Uncertain-aware Reward Model Ensemble (URME) are proposed to model and quantify the uncertainty during reward modeling. Unlike previous methods that deterministically map hidden states to reward scalars, URM and URME model the distribution of rewards and evaluate prediction confidence by uncertainty quantification. Notably, among RMs with 8B or smaller model size, URM achieves state-of-the-art performance on RewardBench, surpassing a number of larger models. Empirical evidence further validates that through uncertainty quantification, URM and URME can effectively evaluate their level of knowledge for input data, leading to more reliable reward predictions.
The limitation of our paper is that our experiment for generation improvements is limited to best-of-$n$ due to the limit of computation resources. In the future, we plan to introduce URM and URME to prevailing LLM alignment methods like RLHF Ouyang et al. (2022) and iterative DPO Yuan et al. (2024). Another direction to look at is model merging for URMs in the weight space (Ramé et al., 2024), which demonstrate competitive efficiency and robustness compared to ensembles. A simple empirical study to RM merging with URMs is included in the appendix C.1.
@@ A.1 Training Details
The gating layer in URM consists of 2 hidden layers, both with 4096 hidden size. The activation function in the gating layer is SELU (Klambauer et al., 2017), which induces self-normalizing properties. We take 4k response pairs from Skywork-reward-preference-80k (Liu & Zeng, 2024) as the validation set. The other data is used for training. We train the gating layer with batchsize 256 for 4000 steps.

### v2

@@ 4.2 Uncertainty-aware Reward Model Ensemble
MODEL BASE SCORE CHAT C-HARD SAFETY REASON
URM(S) Llama3.1-8B 92.9 95.5 88.2 91.1 97.0
Skywork-RM Llama3.1-8B 92.5 95.8 87.3 90.8 96.2
URM(F) Llama3-8B 89.9 96.9 78.7 88.2 95.7
Fsfairx-RM Llama3-8B 84.4 99.4 65.1 86.8 86.4
SFR-Judge-r Llama3.1-70B 92.7 96.9 84.8 91.6 97.6
Nemotron-RM Nemotron4-340B 92.0 95.8 87.1 91.5 93.6
GRM Llama3-8B 91.5 95.5 86.2 90.8 93.6
ArmoRM Llama3-8B 90.4 96.9 76.8 90.5 97.3
InternLM2-RM InternLM2-20B 90.2 98.9 76.5 89.5 95.8
SteerLM-RM Llama3-70B 88.8 91.3 80.3 92.8 90.6
GPT-4o - 86.7 96.1 76.1 88.1 86.6
Table 1: Results on RewardBench. RewardBench evaluates four abilities: Chat, Chat-Hard (C-HARD), Safety and Reasoning. URM(S) and URM(F) are trained from base model Skywork-RM and Fsfairx-RM respectively.
Bootstrap ensemble of models is simple and effective for epistemic uncertainty quantification compared with other methods (Neal, 2012; Hernández-Lobato and Adams, 2015; Blundell et al., 2015).
@@
Accurate reward prediction is crucial in LLM alignment, as it fundamentally steers the learning process. Thus, we can adopt a filtering strategy to discard or penalize data with highly uncertain reward predictions, since RMs may exhibit poor generalization and lack sufficient knowledge to provide reliable feedbacks for them. In this way, we can prevent LLMs from learning undesired behaviors, promoting a more controlled and trustworthy alignment process.
@@ 5.1 Experiment Settings
In our experiment, URM is based on Llama3 and Llama3.1 with 8 billion parameters. Before adding the probabilistic value head, we initialize URM’s base model with weights from Liu and Zeng (2024) for Llama3.1 URM and Dong et al. (2023a) for Llama3 URM. More information on URM training and implementation is given in the appendix A.1, C.1. URME have 3 URMs with different random seeds, probabilistic value head initialization and mini-batches of training data.
We utilize HelpSteer 2 (Wang et al., 2024c) as the training dataset to train the base model and the probabilistic value head for 1 epoch with learning rate $2\times 10^{-6}$.
RewardBench (Lambert et al., 2024) , our evaluation benchmark for RMs, has 2985 questions and response pairs. For multi-attribute RMs and BT-model RMs, a prediction for a response pair is correct if the RM gives a higher reward to the chosen response than the rejected response. For generative models, RewardBench evaluates them via LLM-as-a-judge (Zheng et al., 2023). If the generative model prioritizes the chosen response than the rejected response, the prediction is seen as correct. To test URM and URME’s ability in improving LLMs’ generation quality, we evaluate URM and URME with best-of-$n$ sampling (Stiennon et al., 2020) on AlpacaEval (Li et al., 2023). To evaluate URM’s efficacy in improving LLM alignment, we apply URMs to iterative DPO (Xiong et al., 2024) and RLHF (Ouyang et al., 2022) and compare the improvements over the SFT model.
@@ 5.2.1 Overall Results
Table 1 gives the results on RewardBench. The compared baselines include multi-attribute RMs (Nemotron4-Reward (Adler et al., 2024), ArmoRM (Wang et al., 2024a), SteerLM-RM (Dong et al., 2023b)), BT-model RMs (Skywork-RM (Liu and Zeng, 2024), GRM (Yang et al., 2024), InternLM2-RM (Cai et al., 2024)) and generative RMs (SFR-Judge-r and GPT-4o (OpenAI, 2024)). SFR-Judge-r is a chatbot developed by Salesforce based on Llama3.1-70B.
The results on RewardBench confirm URM’s strong ability in reward modeling. URM(S) and URM(F) both outperform their base model Skywork-RM and Fsfairx-RM. URM also defeats a number of larger-scale models. Especially, compared to Nemotron-RM and ArmoRM which are also a multi-attribute RMs, URM’s better performance shows the efficacy of modeling human preferences as distributions.
Figure 3: (a) Attribute score distributions modeled by URM. Means and variances are estimated and averaged by OOD and in-distributional samples separately. (b-c) URME’s uncertainty estimation on OOD and in-distributional samples. (d) URM and URME’s evaluation accuracy when uncertainty is within the threshold. The results confirm that uncertainty of URM and URME is able to indicate reliability of reward predictions.
@@ 5.2.2 Ablation Study
Here we study the effect of the uncertain-aware value head and different training methods of URMs. To test the applicability of URM, we initialize URM with two different base models: Skywork-RM (Liu and Zeng, 2024) and Fsfairx-RM (Dong et al., 2023a). Table 2 gives the results of our ablation study. ’URM-Det’ refers to the model with a deterministic value head to directly map hidden states to score values instead of sampling in URM. All other components of URM-Det are kept the same as URM. URM-Reg is an URM trained with the attribute regression loss function in Eq. 3, while URM-MLE is trained via maximum likelihood estimation. Since the dataset Helpsteer 2 for our attribute prediction has already been used in the base model Skywork-RM, URM-Det and URM-MLE do not demonstrate improvement over the base model, and only URM-Reg surpasses the base model by modeling the preference distributions. But with base model Fsfairx-RM not trained with Helpsteer 2 previously, all our models demonstrate significant improvement over the base model. Especially, URM trained via attribute regression significantly outperform its counterpart with MLE loss. However, although URM-Reg has better performance in prioritizing chosen responses over the rejected, URM-MLE demonstrates better uncertainty quantification and distribution modeling ability. We theoretically illustrate this phenomenon in the appendix B. Thus, for other studies involving uncertainty quantification, we use URMs trained via the MLE loss.
Skywork-RM Fsfairx-RM
Base 92.5 84.4
URM-Det 92.0±$\pm 0.1$ 88.3 ±$\pm 0.3$
URM-Reg 92.9±$\pm 0.1$ 89.9±$\pm 0.2$
URM-MLE 91.7±$\pm 0.3$ 87.6 ±$\pm 0.4$
Table 2: Alation study results on two different base models. URM-Det is the ablation replacing the uncertainty-aware value head with a deterministic value head.
Table 2 indicates regression-based training methods achieve higher scores on RewardBench. This could potentially be credited to the high quality of Helpsteer 2 dataset, which is meticulously processed. This quality enables even the simplest direct attribute regression to deliver substantial performance improvements, as shown by URM-Det against base model Fsfairx-RM. However, the introduction of noise via the sampling-based scores in URM-Reg makes URMs more robust in distinguishing between chosen and rejected responses and outperforms URM-Det. Despite this, we anticipate that URM-MLE would prove more successful on real-world datasets, which often encompass a wide spectrum of data quality, so that modeling distributions of the scores becomes more necessary.
@@ 5.2.3 Uncertainty Quantification
Figure 4: Results of BoN, iterative DPO and RLHF. (b) Results of iterative DPO. Since URM is not specifically trained to match the initial SFT model, it produces numerous OOD evaluations and performs poorly. However, by leveraging uncertainty to identify and filter out data with OOD evaluations, URM even outperforms the recommended RM, which closely matches the SFT model. (c) Results of RLHF with PPO. Threshold 0 means no training (as no data is below the threshold), and threshold ’inf’ means keeping all rewards regardless of uncertainty. Restricting LLMs from generating texts with high reward prediction uncertainty significantly enhances performance.
@@
Now we study the uncertainty quantification of URM and URME and the relationship between uncertainty and reward reliability. Given the challenge inherent in identifying what precisely is OOD for LLMs, we adopt numeric calculations as simulated OOD data. This is because LLMs are known to underperform in this skill area. Details are given in the appendix A.2.
@@
Fig. 3(a) gives the attribution score attributions of OOD and in-distribution data modeled by URM. Due to the lack of knowledge to accurately evaluate the OOD data, the modeled distributions for OOD data have significantly larger variance and are much closer to uniformity than for in-distributional data. Therefore, this uncertainty quantified by the variance can serve as an informative tool for identifying and filtering out OOD data, where reward models exhibit a tendency towards making uniform guess than providing an accurate evaluation. This strategy ensures the evaluated outcomes are both more dependable and robust.
@@
We quantify uncertainty in URME with two metrics: maximum reward gaps in Eq. 4 and maximum covariance norms in Eq. 5. URME uncertainty quantification results are given in Fig. 3(b), (c). Quantified uncertainty under two metrics both indicate that URME is substantially more uncertain on OOD data. The results confirm that when the URMs lack relevant knowledge to make accurate reward predictions, they will diverge with each other, demonstrating significant discrepancies.
@@
To test whether quantifying uncertainty is able to improve reliability of reward predictions, we utilize 2k prompts from the held-out validation set and evaluate their rewards and uncertainties. Fig. 3(d) gives URM and URME’s evaluation accuracy with different uncertainty threshold. In this setup, prompts and responses (either the chosen or rejected) with uncertainty larger than the threshold are filtered out. URME uses reward gaps to quantify uncertainty and URM’s uncertainty is quantified by summation of attribute distributions’ variance. The results validate our claim that reward predictions with low uncertainty are more reliable than those with high uncertainty. Therefore, through uncertainty quantification, we can identify unreliable reward predictions and filter them out, leading to improved rewards reliability.
5.2.4 Generations and Alignment Improvement
To test URM and URME’s ability in improving LLMs’ generation and alignment, we conduct experiments on AlpacaEval Li et al. (2023) with Best-of-n (BoN) sampling, iterative DPO (Xiong et al., 2024) and RLHF (Ouyang et al., 2022).
Settings For BoN experiments, we use URM and URME to select best responses to compare against the reference answers. Our experiments of iterative DPO and RLHF follows the implementation of OpenRLHF (Hu et al., 2024). All settings are set as default (e.g., recommended SFT model and dataset) unless stated otherwise. For iterative DPO, we start with an SFT model as the initial point. In each iteration, the model generates multiple responses for 8,000 training prompts. These responses are then ranked by RMs to create chosen-rejected pairs, which are subsequently used for DPO training. In experiments with URM, we only keep $50\%$ of the training pairs with lower uncertainty to ensure reliable evaluations. For RLHF, we apply PPO to the SFT model. In URM experiments, we penalize responses with uncertainty higher than a threshold, so that the model can keep away from generating texts with unreliable and inaccurate reward predictions. To compare the performance of different models, we calculate the win rate using LLM-as-a-Judege with GPT-4-0125-preview as the judge. To eliminate positional bias of LLM judges, we switch positions of models and evaluate twice to ensure fairness of evaluation results. More details are given in the appendix A.3. For BoN, we evaluate Llama3-8B-Instruct’s win rate against the reference answers as $n$ increases, while for iterative DPO and RLHF experiments, we compare the models’ performance against the SFT model.
Results Fig. 4(a) demonstrates that as the number of samples increases, both URM and URME effectively evaluate response quality, improving the generative performance of the baseline model.
Fig. 4(b) gives iterative DPO results. URM is not specifically trained to match the SFT model and may generate numerous OOD evaluations. Without uncertainty-based filtering, URM performs significantly worse due, even leading to model degeneration. In contrast, applying uncertainty-based filtering allows URM to mitigate OOD effects by discarding data with unreliable rewards, ultimately enabling it to outperform the recommended RM, which is trained to closely match the SFT model.
Fig. 4(c) depicts the RLHF results. When high-uncertainty responses are not penalized (threshold ’inf’), the win rate against the SFT model remains at $60.2\%$, due to unreliable reward predictions—similar to the behavior observed in iterative DPO. However, penalizing high-uncertainty responses leads to significant performance gains (threshold 10, 30, 50). Notably, the PPO model achieves its highest performance at a threshold of 50. For lower thresholds 10 and 30, fewer responses exhibit acceptable uncertainty, limiting the number of useful data to align the LLM effectively.
6 Conclusions
In this paper, we introduced URM and URME. By incorporating uncertainty into reward modeling, URM and URME enhance the reliability of reward predictions, effectively identifying instances where predictions are less trustworthy. This capability significantly mitigates the risks of propagating unreliable evaluations during LLM alignment. Empirical results demonstrates the proposed models’ effectiveness and reliability in reward prediction. The integration of URM into LLM fine-tuning, such as iterative DPO and RLHF, highlights the practical utility of our approach in improving alignment and generation quality of LLMs.
@@ A.1 Training Details
To demonstrate the wide applicability of URM, we implement URM(S) with a gating layer while attribute scores in URM(F) are combined via pre-determined weights from Adler et al. (2024). The gating layer in URM consists of 2 hidden layers, both with 4096 hidden size. The activation function in the gating layer is SELU (Klambauer et al., 2017), which induces self-normalizing properties. After obtaining the attribute-specific uncertain-aware probabilistic value head and base model, we keep them frozen and train the gating layer on Skywork-reward-preference-80k (Liu and Zeng, 2024) for 4000 steps with batchsize 256. We train the gating layer with batchsize 256 for 4000 steps. During training, we held out 4k data from the dataset as validation set to choose the checkpoint with highest validation accuracy.
Artifacts used in this paper include OpenRLHF (Hu et al., 2024), RLHF-Reward-Modeling (Dong et al., 2024), HelpSteer 2 (Wang et al., 2024c), Skywork-reward-preference-80k (Liu and Zeng, 2024) and AlpacaEval (Li et al., 2023), whose licenses all permit academic usage.
Recently, RewardBench found that there is overlap between the benchmark’s test set and training set Skywork-reward-preference-80k-v0.1, which leads to unintentional contamination. But, the gating layer in URM is not a necessary component. By removing the gating layer, the performance of URM is {’Chat’: 0.955, ’Chat Hard’: 0.864, ’Safety’: 0.909, ’Reasoning’: 0.976}, not much difference with using a gating layer. And we do not use the gating layer in URM(F) and later alignment experiment with URMs.